OpenStudy (rainbow_dashie):

Kate is making hamburgers for her friends and has 13 1/2 pounds of hamburger. How many 1/2-pound hamburgers can she make?

OpenStudy (rainbow_dashie):

@RyleeC @amistre64

OpenStudy (anonymous):

Divide 13 1/2 by 1/2

OpenStudy (anonymous):

Or 13.5 / 0.5

OpenStudy (kropot72):

This can be solved by using the decimal values of the given numbers. \[\large Number\ of\ burgers=\frac{13.5}{0.5}\]

OpenStudy (anonymous):

13 (1/2) / 1/2 = (27/2)/1/2 = 27/4 = ans